Suppose a jar contains 7 red marbles and 32 blue marbles. If you reach in the jar and pull out one marble at random, do not replace the marble, and reach in the jar to pull out a second marble. Find the probability that both are red. Write your answer as a reduced fraction
Answer #1

Total number of marbles = 7+32 = 39

Number of red marbles = 7

Probability that the first marble is red = 7/39

After picking the first red marble, there are 6 red marbles left in the jar and a total of (6+32) = 38 marbles left in the jar.

Probability that the second marble is red = 6/38

Hence, the probability that both marbles are red = (7/39)×(6/38) = 7/247